# Insect Identification in the Wild: The AMI Dataset

### v1

@@ Abstract
Insects represent half of all global biodiversity, yet many of the world’s insects are disappearing, with severe implications for ecosystems and agriculture. Despite this crisis, data on insect diversity and abundance remain woefully inadequate, due to the scarcity of human experts and the lack of scalable tools for monitoring. Ecologists have started to adopt camera traps to record and study insects, and have proposed computer vision algorithms as an answer for scalable data processing. However, insect monitoring in the wild poses unique challenges that have not yet been addressed within computer vision, including the combination of long-tailed data, extremely similar classes, and significant distribution shifts. We provide the first large-scale machine learning benchmarks for fine-grained insect recognition, designed to match real-world tasks faced by ecologists. Our contributions include a curated dataset of images from citizen science platforms and museums, and an expert-annotated dataset drawn from automated camera traps across multiple continents, designed to test out-of-distribution generalization under field conditions. We train and evaluate a variety of baseline algorithms and introduce a combination of data augmentation techniques that enhance generalization across geographies and hardware setups. Code and datasets are made publicly available [^1] https://github.com/RolnickLab/ami-dataset.
@@ 2 Related Work
Insect Camera Traps. The use of camera traps for studying insect biodiversity is relatively recent. Geissmann et al. introduced the Sticky Pi device, which uses sticky cards to capture (and kill) insects, which are then automatically photographed [32]. Bjerge et al. introduced an automated camera trap using UV light to attract and photograph moths and other insects (without killing them) [21], which was built on by subsequent works [20, 18]. A similar approach focused on diurnal pollinators via a camera trap installed above flowers [17, 19]. These works also propose deep learning methods for processing the resulting images, considering both object detection and classification of a small number of insect categories. The lack of fine-grained classification is largely due to the time required to annotate images from the traps, making it difficult to obtain a large enough labeled dataset from trap data alone.
@@ 4 The AMI Dataset
In this section, we describe the construction of the datasets that we release in this work: AMI-Traps and AMI-GBIF. For AMI-Traps, we collected insect camera trap data by field deployments in three regions: NE-America (Northeast of North America), which includes data from Quebec in Canada and Vermont in the USA; W-Europe (Western Europe), which includes data from the UK and Denmark; and C-America (Central America), which includes data from Panama. For AMI-GBIF, we compiled public data for the same three regions using the Global Biodiversity Information Facility (GBIF) [7]. The GBIF platform provides access to biodiversity data aggregated from multiple sources, such as museums, research institutions, and citizen science platforms such as iNaturalist [4], Observation.org [5], and BugGuide [1]. For brevity, we sometimes refer to AMI-Traps and AMI-GBIF as AMI-T and AMI-G, respectively. Figure 3 shows examples of images from the two datasets and Tab. 1 compares the number of images and coverage of classes at different taxonomic levels.
@@ 4.2 AMI-GBIF Dataset
Statistics. The final dataset contains $\sim$2.5M images from 5,364 species, 1,734 genera and 77 families of moths, along with 350,000 additional images of non-moth groups. The binary dataset contains 700,000 images, composed of 350,000 moth images and the 350,000 non-moth images. The fine-grained dataset contains 2,564,392 moth images, and Tab. 1 shows the number of images and species for each region. Though species are selected based on their presence in regions of interest, we include all GBIF observations of these species from anywhere on Earth (see Fig. 0.A.6 for the global distribution of images). As in other biodiversity datasets, the number of images per species in AMI-GBIF follows a long-tailed distribution (see Fig. 0.C.7). Specifically, only 3,099 species have more than 100 images and are considered under the “many-shot” bucket, 1,358 species contains between 20 and 100 images and are considered “medium-shot”, and there are 907 species with less than 20 training images, considered as “few-shot” categories.
@@ 5 Methods
In this section, we describe the methods used to assess the difficulty of the proposed benchmark tasks. We evaluate a variety of high-performing image classification architectures and use extensive data augmentation to address the challenges of domain adaptation. Motivated by the particular structure of the AMI-Traps dataset, we add a simple mixed-resolution (MixRes) data augmentation technique. As we describe in the following section, our approaches perform quite well, with MixRes leading to an especially significant performance improvement. However, more complex approaches for domain adaptation may further improve the performance on AMI-Traps in particular.
@@
Data Augmentation Approaches. One characteristic of the in-the-wild images is the large variety of crops due to the diversity of insect sizes. A considerable number of them have small resolutions, below 64 $\times$ 64. At this scale, some important visual features that the model may have learned from high-resolution GBIF images might not be identifiable, leading to inaccurate predictions. This issue can be further amplified when these low-resolution crops are scaled up to the input model size. To compel the model to learn features for low-resolution inputs and avoid bias towards visual features only distinguishable in high-resolution images from other domains, we use a simple data augmentation technique we refer to as Mixed Resolution (MixRes) that simulates this low-resolution input during training (see Fig. 0.D.8 in supplementary). This approach consists of randomly resizing the training input image to 25% or 50% of the model input size with a probability $\rho$ for each case. In our experiments $\rho$ is set to $0.25$, maintaining the original input size with probability $0.5$. We employ MixRes in all experiments unless otherwise stated.
In addition to MixRes, we also use a set of standard data augmentation operations: random resize crop with scale of 0.3 to 1.0 and an aspect ratio between 3/4 and 4/3, random horizontal flip with 50% probability, and RandAugment [24] with $N=2$ operations and magnitude $M=9$. Finally, the image is normalized using the mean and std of ImageNet and scaled to the input network size.
Implementation Details. We train all models with ImageNet-pretrained weights from the timm library [78] (for ConvNext and ViT we used ImageNet-22K weights), using only the AMI-GBIF training sets. Each model is trained with a batch size of 128 and the AdamW optimizer [49] for 30 epochs. The initial learning rate is set to 0.001, with a weight decay of 1e-5. The learning rate is linearly warmed up from 0 to the initial value for 2 epochs and then decayed using cosine scheduling [36]. We also apply label smoothing [67] with a value of 0.1.
Inference Settings. At test time, the insect crops from AMI-Traps are padded with black pixels to make them square, preventing the distortion of the shape of the insect before prediction. Next, the images are normalized with the mean and std of ImageNet and, finally, resized to the model input size. No padding is done for the test images from AMI-GBIF.
@@
Evaluation Metrics. During the binary classification step, it is important to achieve high precision by filtering out as many non-moths as possible to prevent the propagation of errors in the pipeline. At the same time, it is also important to assess recall because if too many moths are discarded at this step, it may impact the utility of the data for ecological studies. Therefore, for the binary classification task, along with accuracy, we also consider precision, recall, and F1-score as evaluation metrics. For the fine-grained classification task, we evaluate the model in terms of accuracy at the species, genus and family levels. We present the results averaged over three independent runs.
@@ 6 Experiments and Results
Binary Classification Results. This task involves distinguishing moths from non-moths. Here we train a ResNet50 model on the AMI-GBIF dataset and evaluate performance both on the AMI-GBIF held-out test set and on the entirety of the AMI-Traps dataset. On AMI-GBIF (in-distribution), the model achieves a test accuracy of $97.77\%\pm 0.02\%$. Table 2 shows the results for the AMI-Traps dataset in more detail. We observe that the model has a high recall but a lower precision, which means it is over-predicting the moth class. Also, the bigger the image crop, the better the accuracy of the insect classification; many non-moths where the model errs are extremely small and are therefore highly blurred in the trap images. Table 0.E.7 illustrates the improvement in model performance when crops are square padded (our default at inference time) versus an ablation where padding is not performed.
@@
Fine-grained Classification Results. The fine-grained classification task aims to identify moths to the species level (and secondarily to genus or family level). In our experiments, we train models on AMI-GBIF data for each of the three regions of the AMI dataset, evaluating both on AMI-GBIF (in-distribution) and AMI-Traps (out-of-distribution) data (see Tab. 3). Not all moth species known to exist in a region appear on camera traps; hence, the list of species in AMI-Traps is a subset of species in AMI-GBIF for a given area (see Tab. 1). Therefore, we also evaluate the models on an in-distribution subset (Test‡) from AMI-GBIF containing only the species present in AMI-Traps. Meanwhile, for AMI-Traps, there are 16 species not present in AMI-GBIF that were removed during our evaluation procedure (these represent unexpected species not included in the regional checklists used to define AMI-GBIF). Overall, the best model at the species level is ConvNeXt-B for both AMI-GBIF and AMI-traps. For C-America, annotations are mostly provided at the genus or family level, due to the challenging nature of visual identification of moths in this biodiversity-rich region. Furthermore, while the genus and family accuracies are similar between N-America and W-Europe, they are much lower for C-America. This can be attributed to the fact that there are 10 times fewer images in the training set for this region. Note that for genus and family evaluation, there are some cases where the accuracy is lower than the species level. This occurs because those higher taxonomic levels include crops that have labels only at higher levels. Additional results for long-tail accuracies for AMI-GBIF are available in Table 0.E.8 in the supplementary materials.
One important point for practical use cases is how accuracy changes if predictions are only accepted above a certain “confidence threshold”. Thresholding predictions can increase the reliability of model outputs for downstream applications, though it comes with a trade-off of rejecting a certain fraction of predictions. Fig. 4 shows the curves of the accuracy and rejection rate as functions of the confidence threshold for ResNet50 models.
@@
Test Test‡ Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
NE-America ResNet50 87.56$\pm$0.12 89.39$\pm$0.14 71.86$\pm$1.02 91.10$\pm$0.70 80.07$\pm$0.49 93.79$\pm$0.24 90.44$\pm$0.41 98.17$\pm$0.09
MBNetV3L 83.73$\pm$0.17 85.80$\pm$0.29 67.41$\pm$0.72 87.46$\pm$1.05 74.83$\pm$1.30 92.27$\pm$1.06 89.31$\pm$0.62 97.96$\pm$0.06
ConvNeXt-T 89.81$\pm$0.05 91.21$\pm$0.06 74.89$\pm$0.05 91.92$\pm$0.20 83.26$\pm$1.22 93.80$\pm$0.62 90.61$\pm$0.53 98.09$\pm$0.03
ConvNeXt-B 90.45$\pm$0.03 91.66$\pm$0.03 76.75$\pm$1.52 93.51$\pm$0.59 85.46$\pm$0.83 94.40$\pm$0.64 91.07$\pm$0.29 98.18$\pm$0.11
ViT-B/16 86.39$\pm$0.05 87.94$\pm$0.10 77.43$\pm$0.79 93.66$\pm$0.74 85.57$\pm$0.98 95.34$\pm$0.04 91.45$\pm$0.34 98.31$\pm$0.08
W-Europe ResNet50 86.45$\pm$0.08 90.25$\pm$0.03 79.39$\pm$0.39 91.43$\pm$0.23 78.95$\pm$0.26 91.31$\pm$0.30 89.13$\pm$0.42 96.98$\pm$0.39
MBNetV3L 82.20$\pm$0.05 86.47$\pm$0.13 76.30$\pm$1.29 89.80$\pm$0.52 76.35$\pm$1.19 90.18$\pm$0.39 88.33$\pm$0.47 97.05$\pm$0.26
ConvNeXt-T 88.82$\pm$0.03 92.05$\pm$0.05 82.44$\pm$0.89 92.28$\pm$0.62 81.86$\pm$0.79 92.15$\pm$0.34 90.08$\pm$0.59 97.57$\pm$0.24
ConvNeXt-B 89.35$\pm$0.01 92.48$\pm$0.05 82.87$\pm$0.39 92.90$\pm$0.08 82.45$\pm$0.44 92.84$\pm$0.18 91.27$\pm$0.60 97.81$\pm$0.10
ViT-B/16 85.45$\pm$0.03 89.22$\pm$0.11 82.44$\pm$0.36 92.88$\pm$0.06 83.01$\pm$0.26 92.65$\pm$0.18 91.68$\pm$0.17 97.70$\pm$0.15
C-America ResNet50 87.46$\pm$0.23 95.92$\pm$0.68 - - 47.44$\pm$1.11 65.38$\pm$0.00 69.18$\pm$0.24 87.59$\pm$0.24
MBNetV3L 83.68$\pm$0.24 92.89$\pm$0.27 - - 41.03$\pm$2.22 62.18$\pm$4.44 68.34$\pm$1.69 86.89$\pm$0.64
ConvNeXt-T 87.65$\pm$0.11 95.46$\pm$0.46 - - 55.77$\pm$3.33 66.03$\pm$4.00 69.46$\pm$0.42 88.56$\pm$1.05
ConvNeXt-B 88.62$\pm$0.09 96.22$\pm$0.26 - - 51.92$\pm$3.85 67.31$\pm$1.93 72.24$\pm$1.47 89.54$\pm$0.84
ViT-B/16 79.80$\pm$0.26 88.36$\pm$1.02 - - 47.44$\pm$4.44 65.39$\pm$3.33 69.87$\pm$1.51 85.49$\pm$1.28
@@
Table 4: Ablation study on data augmentation techniques using a ResNet50 for the NE-America region evaluated on top-1 species accuracy. “RandAug” denotes the combination of the RandAugment and random resize crop data augmentation techniques.
Resolution RandAug MixRes AMI-GBIF AMI-Traps
@@
85.14$\pm$0.08 58.41$\pm$0.30
128 87.56$\pm$0.12 71.86$\pm$1.02
224 91.51$\pm$0.11 68.41$\pm$1.06
Figure 4: The left y-axis compares the top-1 accuracy as a function of model (ResNet50) confidence, while the right y-axis shows the rejection rate for the corresponding thresholds for the NE-America (left) and W-Europe regions (right). At a confidence of 50%, the model has over 90% accuracy at all taxonomic levels. This analysis is skipped for the C-America which has sparse taxonomic coverage in the AMI-T dataset.

### v2

@@ Abstract
Insects represent half of all global biodiversity, yet many of the world’s insects are disappearing, with severe implications for ecosystems and agriculture. Despite this crisis, data on insect diversity and abundance remain woefully inadequate, due to the scarcity of human experts and the lack of scalable tools for monitoring. Ecologists have started to adopt camera traps to record and study insects, and have proposed computer vision algorithms as an answer for scalable data processing. However, insect monitoring in the wild poses unique challenges that have not yet been addressed within computer vision, including the combination of long-tailed data, extremely similar classes, and significant distribution shifts. We provide the first large-scale machine learning benchmarks for fine-grained insect recognition, designed to match real-world tasks faced by ecologists. Our contributions include a curated dataset of images from citizen science platforms and museums, and an expert-annotated dataset drawn from automated camera traps across multiple continents, designed to test out-of-distribution generalization under field conditions. We train and evaluate a variety of baseline algorithms and introduce a combination of data augmentation techniques that enhance generalization across geographies and hardware setups. The dataset is made publicly available [^1] https://github.com/RolnickLab/ami-dataset.
@@ 2 Related Work
Insect Camera Traps. The use of camera traps for studying insect biodiversity is relatively recent. Geissmann et al. [32] introduced the Sticky Pi device, which uses sticky cards to capture (and kill) insects, which are then automatically photographed. Bjerge et al. [21] introduced an automated camera trap using UV light to attract and photograph moths and other insects (without killing them), which was built on by subsequent works [20, 18]. A similar approach focused on diurnal pollinators via a camera trap installed above flowers [17, 19]. These works also propose deep learning methods for processing the resulting images, considering both object detection and classification of a small number of insect categories. The lack of fine-grained classification is largely due to the time required to annotate images from the traps, making it difficult to obtain a large enough labeled dataset from trap data alone.
@@ 4 The AMI Dataset
In this section, we describe the construction of the datasets that we release in this work: AMI-Traps and AMI-GBIF. For AMI-Traps, we collected insect camera trap data by field deployments in three regions: NE-America (Northeast of North America), which includes data from Quebec in Canada and Vermont in the USA; W-Europe (Western Europe), which includes data from the UK and Denmark; and C-America (Central America), comprising data from Panama. For AMI-GBIF, we compiled public data for the same three regions using the Global Biodiversity Information Facility (GBIF) [7]. The GBIF platform provides access to biodiversity data aggregated from multiple sources, such as museums, research institutions, and citizen science platforms such as iNaturalist [4], Observation.org [5], and BugGuide [1]. For brevity, we sometimes refer to AMI-Traps and AMI-GBIF as AMI-T and AMI-G, respectively. Figure 3 shows examples of images from the two datasets and Tab. 1 compares the number of images and coverage of classes at different taxonomic levels.
@@ 4.2 AMI-GBIF Dataset
Statistics. The final dataset contains $\sim$2.5M images from 5,364 species, 1,734 genera and 77 families of moths, along with 350,000 additional images of non-moth groups. The binary dataset contains 700,000 images, composed of 350,000 moth images and the 350,000 non-moth images. The fine-grained dataset contains 2,564,392 moth images, and Tab. 1 shows the number of images and species for each region. Though species are selected based on their presence in regions of interest, we include all GBIF observations of these species from anywhere on Earth (see Fig. 0.A.5 for the global distribution of images). As in other biodiversity datasets, the number of images per species in AMI-GBIF follows a long-tailed distribution (see Fig. 0.C.6). Specifically, only 3,099 species have more than 100 images and are considered under the “many-shot” bucket, 1,358 species contains between 20 and 100 images and are considered “medium-shot”, and there are 907 species with less than 20 training images, considered as “few-shot” categories.
@@ 5 Methods
In this section, we describe the methods used to assess the difficulty of the proposed benchmark tasks. We evaluate a variety of high-performing image classification architectures and use extensive data augmentation to address the challenges of domain shift between the two datasets. Motivated by the particular structure of the AMI-Traps dataset, we add a mixed-resolution (MixRes) data augmentation technique. As we describe in the following section, our approaches perform quite well, with MixRes leading to an especially significant performance improvement.
@@
Data Augmentation Approaches. One characteristic of the in-the-wild images is the large variety of crops due to the diversity of insect sizes. A considerable number of them have small resolutions, below 64 $\times$ 64. At this scale, some important visual features that the model may have learned from high-resolution GBIF images might not be identifiable, leading to inaccurate predictions. This issue can be further amplified when these low-resolution crops are scaled up to the input model size. To compel the model to learn features for low-resolution inputs and avoid bias towards visual features only distinguishable in high-resolution images from other domains, we use a data augmentation technique we refer to as Mixed Resolution (MixRes) that simulates this low-resolution input during training (see Fig. 0.D.7 in supplementary). This approach consists of randomly resizing the training input image to 25% or 50% of the model input size with a probability $\rho$ for each case. In our experiments $\rho$ is set to $0.25$, maintaining the original input size with probability $0.5$. We employ MixRes in all experiments unless otherwise stated.
In addition to MixRes, we also use a set of standard data augmentation operations: random resize crop with scale of 0.3 to 1.0 and an aspect ratio between 3/4 and 4/3, random horizontal flip with 50% probability, and RandAugment [24] with $N=2$ operations and magnitude $M=9$. Finally, the image is normalized using the mean and standard deviation of ImageNet and scaled to the input network size.
Implementation Details. We train all models with the AMI-GBIF training sets using the ImageNet-1K pretrained weights from the timm library [78] (for ConvNext and ViT we used ImageNet-22K weights). Each model is trained with a batch size of 128 and the AdamW optimizer [49] for 30 epochs. The initial learning rate is set to 0.001, with a weight decay of 1e-5. The learning rate is linearly warmed up from 0 to the initial value for 2 epochs and then decayed using cosine scheduling [36]. We also apply label smoothing [67] with a value of 0.1.
Inference Settings. At test time, the insect crops from AMI-Traps are padded with black pixels to make them square, preventing the distortion of the shape of the insect before prediction. Next, the images are normalized with the mean and standard deviation of ImageNet and, finally, resized to the model input size. No padding is done for the test images from AMI-GBIF.
@@
Evaluation Metrics. During the binary classification step, it is important to achieve high precision by filtering out as many non-moths as possible to prevent the propagation of errors in the pipeline. At the same time, it is also important to assess recall because if too many moths are discarded at this step, it may impact the utility of the data for ecological studies. Therefore, for the binary classification task, along with accuracy, we also consider precision, recall, and F1-score as evaluation metrics. For the fine-grained classification task, we evaluate the model in terms of micro- and macro-averaged accuracy at the species, genus and family levels. Micro-averaging aggregates all class samples to calculate the accuracy while the macro-average is the average of per-class accuracy. Unless otherwise stated, the reported accuracy is micro-averaged. We conduct each experiment for three independent runs and report the mean and standard deviation.
@@ 6 Experiments and Results
Binary Classification Results. This task involves distinguishing moths from non-moths. Here we train a ResNet50 model on the AMI-GBIF dataset and evaluate performance both on the AMI-GBIF held-out test set and on the entirety of the AMI-Traps dataset. On AMI-GBIF (in-distribution), the model achieves a test accuracy of $97.77\%\pm 0.02\%$. Table 2 shows the results for the AMI-Traps dataset in more detail. We observe that the model has a high recall but a lower precision, which means it is over-predicting the moth class. Also, the bigger the image crop, the better the accuracy of the insect classification; many non-moths where the model errs are extremely small and are therefore highly blurred in the trap images. Table 0.E.7 in the supplementary illustrates the improvement in model performance when crops are square padded (our default at inference time) versus an ablation where padding is not performed.
@@
Fine-grained Classification Results. The fine-grained classification task aims to identify moths to the species level (and secondarily to genus or family level). In our experiments, we train models on AMI-GBIF data for each of the three regions of the AMI dataset, evaluating both on AMI-GBIF (in-distribution) and AMI-Traps (out-of-distribution) data (see Tab. 3). Not all moth species present in a region appear regularly at camera traps; hence, the list of species in AMI-Traps is a subset of species in AMI-GBIF for a given area (see Tab. 1). Therefore, we also evaluate the models on an in-distribution subset (Test‡) from AMI-GBIF containing only the species present in AMI-Traps. Meanwhile, for AMI-Traps, there are 16 species not present in AMI-GBIF that were removed during our evaluation procedure (these represent unexpected species not included in the regional checklists used to define AMI-GBIF). Overall, the best model at the species level is ConvNeXt-B for both AMI-GBIF and AMI-traps. For C-America, annotations are mostly provided at the genus or family level, due to the challenging nature of visual identification of moths in this biodiversity-rich region. Furthermore, while the genus and family accuracies are similar between N-America and W-Europe, they are much lower for C-America. This can be attributed to the fact that there are 10 times fewer images in the training set for this region. Note that for genus and family evaluation, there are some cases where the accuracy is lower than the species level. This occurs because those higher taxonomic levels include crops that have labels only at higher levels. In the supplementary materials, we include additional results for AMI-GBIF long-tail accuracies (Tab. 0.E.8), a pre-training ablation study (Tab. 0.E.9), and a study analyzing classification accuracy as a function of prediction confidence (Fig. 0.E.8).
@@
[…]
NE-America ResNet50 87.56$\pm$0.12 89.39$\pm$0.14 71.86$\pm$1.02 64.96$\pm$2.00 80.07$\pm$0.49 74.66$\pm$0.35 90.44$\pm$0.41 69.30$\pm$4.15
MBNetV3L 83.73$\pm$0.17 85.80$\pm$0.29 67.41$\pm$0.72 61.21$\pm$1.20 74.83$\pm$1.30 71.20$\pm$2.47 89.31$\pm$0.62 64.48$\pm$1.87
ConvNeXt-T 89.81$\pm$0.05 91.21$\pm$0.06 74.89$\pm$0.05 68.00$\pm$0.76 83.26$\pm$1.22 77.26$\pm$0.85 90.61$\pm$0.53 74.60$\pm$0.64
ConvNeXt-B 90.45$\pm$0.03 91.66$\pm$0.03 76.75$\pm$1.52 68.56$\pm$0.62 85.46$\pm$0.83 78.28$\pm$0.14 91.07$\pm$0.29 76.87$\pm$1.10
ViT-B/16 86.39$\pm$0.05 87.94$\pm$0.10 77.43$\pm$0.79 68.02$\pm$0.34 85.57$\pm$0.98 78.97$\pm$0.72 91.45$\pm$0.34 73.66$\pm$1.49
W-Europe ResNet50 86.45$\pm$0.08 90.25$\pm$0.03 79.39$\pm$0.39 72.84$\pm$0.21 78.95$\pm$0.26 73.22$\pm$0.60 89.13$\pm$0.42 75.24$\pm$1.85
MBNetV3L 82.20$\pm$0.05 86.47$\pm$0.13 76.30$\pm$1.29 70.35$\pm$0.93 76.35$\pm$1.19 72.88$\pm$1.28 88.33$\pm$0.47 72.21$\pm$0.38
ConvNeXt-T 88.82$\pm$0.03 92.05$\pm$0.05 82.44$\pm$0.89 74.59$\pm$0.62 81.86$\pm$0.79 74.15$\pm$0.59 90.08$\pm$0.59 74.58$\pm$2.93
ConvNeXt-B 89.35$\pm$0.01 92.48$\pm$0.05 82.87$\pm$0.39 76.41$\pm$0.50 82.45$\pm$0.44 75.98$\pm$1.05 91.27$\pm$0.60 79.39$\pm$2.70
ViT-B/16 85.45$\pm$0.03 89.22$\pm$0.11 82.44$\pm$0.36 75.25$\pm$0.98 83.01$\pm$0.26 75.96$\pm$0.35 91.68$\pm$0.17 75.81$\pm$0.79
C-America ResNet50 87.46$\pm$0.23 95.92$\pm$0.68 - - 47.44$\pm$1.11 66.59$\pm$0.60 69.18$\pm$0.24 42.40$\pm$0.31
MBNetV3L 83.68$\pm$0.24 92.89$\pm$0.27 - - 41.03$\pm$2.22 59.32$\pm$1.18 68.34$\pm$1.69 41.42$\pm$2.98
ConvNeXt-T 87.65$\pm$0.11 95.46$\pm$0.46 - - 55.77$\pm$3.33 70.68$\pm$1.58 69.46$\pm$0.42 45.35$\pm$0.67
ConvNeXt-B 88.62$\pm$0.09 96.22$\pm$0.26 - - 51.92$\pm$3.85 68.63$\pm$1.94 72.24$\pm$1.47 44.31$\pm$1.32
ViT-B/16 79.80$\pm$0.26 88.36$\pm$1.02 - - 47.44$\pm$4.44 64.70$\pm$4.07 69.87$\pm$1.51 42.00$\pm$3.60
@@
[…]
@@
87.41$\pm$0.06 58.41$\pm$0.30
128 89.39$\pm$0.14 71.86$\pm$1.02
224 92.97$\pm$0.17 68.41$\pm$1.06